1. What is the momentum of a golf ball with a mass of 72 g moving at 43 m/s?

2. The impact between the ball and the club lasted for 2.3 x 10³ s. Given the problem in item number 1, what force has acted on the ball? What force has acted on the club?


1
Expert's answer
2021-12-06T09:49:57-0500

Given:

"m=0.072\:\rm kg"

"v=43\:\rm m/s"

"\Delta t=2.3*10^{-3}\rm s"


1. The momentum of the ball

"p=mv=\rm 0.072\:kg*43\: m/s=3.1\: kg*m/s"

2. The force acting on the ball

"F=\frac{\Delta p}{\Delta t}=\rm \frac{3.1\: kg*m/s}{2.3*10^{-3}\: s}=1346\: N"

The force that acted on the club is the same.